OpenStudy (anonymous):

Solve for x. 13(x - 3) = 39 x = 1 x = 4 x = 6 x = 10

OpenStudy (anonymous):

x=6 as 13*3=39

Parth (parthkohli):

13x - 39 = 39 13x = 39 + 39 = 78 x = 78/13 = 6 x = 6

OpenStudy (anonymous):

13x-39=39 13x=39+39 13x=78 x=78/13 x=6